OpenStudy (anonymous):

PLease help. I need help on how to do it not just answer. A gas in a sealed container has a pressure of 133 kPa at a temperature of 30.0°C. If the pressure in the container is increased to 201 kPa, what is the new temperature?

OpenStudy (anonymous):

@ghazi do you think you could help me? please.

OpenStudy (ghazi):

PV=nRT so basically, P is proportional to T that means P=K*T lets say K is some constant so there can be a relation \[\frac{ P _{1}}{ T _{1}}=\frac{ P _{2} }{ T _{2} }\] now put the values and get it done :)

OpenStudy (anonymous):

what would be the value of n and R?

OpenStudy (ghazi):

you dont need that , you just need p and v that you already have

OpenStudy (anonymous):

wait I thought I only had the temperature and the pressure? I'm so sorry, but I'm not understanding how volume goes into that?

OpenStudy (ghazi):

\[\frac{ 133 }{ 30 }=\frac{ 210 }{ T }\] i hope you can do this , and rest volume, moles and gas consant (R) are taken as constant

OpenStudy (anonymous):

47.4 = T?

OpenStudy (ghazi):

correct

OpenStudy (anonymous):

thank you!

OpenStudy (ghazi):

:)